The prism of spectrometer has a refractive angle of 60 degree and is made of glass whoose refractive index for red and violet are 1.514 and 1.530 respectively


Determine the angle of incidence of the light on the prism


1
Expert's answer
2021-05-17T18:35:49-0400

"u_{red}=1.514,u_{violet}=1.530
,
r=60^{\circ}"


Let incidence angle be 'i'-


So Incidence angle-

"\text{sin }i =u_{red}\times sin\dfrac{r}{2}\\[9pt]\text{sin }i=1.514\times sin30=1.514\times 0.5=0.757"


"i=sin^{-1}(0.757)=49.2^{\circ}"


Hence Angle of incidence is "49.2^{\circ}"